# A Biologically Plausible Dense Associative Memory with Exponential Capacity

### v1

@@ 3.2 Experiments
Figure 3a shows the learned basic memories for the MNIST dataset, which correspond to the columns of $\xi$}. As shown in Figure 3b, these basic memories are nearly orthogonal, consistent with Eq. (7).
Figure 3: a) 25 (out of 50) columns of the learned weight matrix, for MNIST images, which serve as basic memories, are shown as two-dimensional images. b) Correlation matrix of the basic patterns, which correspond to the hidden units. c) The network generalizes compositionally, associating unseen cues with interpretable fixed points.
@@
The network learns an effective threshold of $\theta=0.21$ for MNIST and $\theta=0.43$ for CIFAR-10. Note that the statistics of the learned base memories in MNIST and CIFAR-10 differ from one another and from the normal distribution assumed in the theory, which explains the difference between the learned thresholds.
@@
To quantitatively evaluate this behavior, we trained nonlinear classifiers on the recalled representations of the stored images, and tested them on the recalled representations of unseen images (Convolutional neural network (CNN) for visible representations and multilayer perceptron (MLP) for hidden representations). For comparison, we trained CNNs directly on the original stored images and tested them on the original unseen images as well. This allows us to assess the classifiability of the recalled visible representations by the associative memory network relative to the original memories. (see Appendix B for details of the classifiers)
For the MNIST dataset, classification accuracy is high for both the hidden and visible representations. This is a desirable property, as it indicates that the lower-dimensional hidden representations still preserve strong class discriminability. In MNIST, the raw pixel space itself carries strong class structure: two images of the same digit are highly correlated and closer to each other than images of different digits. Consequently, the hidden neurons retain this information almost perfectly, having structured and meaningful encodings in which correlated memories are represented close together and remain classifiable.
For CIFAR-10, classification accuracy is high for the visible representations but low for the hidden ones. This difference arises because, in CIFAR-10, two images of the same class (for example, dogs) are not necessarily correlated in raw pixel space, so the hidden layer, which is a nonlinear transformation of those pixels, does not exhibit a clear class structure. The classifier used for the visible representations in this analysis is a CNN, which, when trained on raw pixel data, first learns a nonlinear transformation that maps images of the same class close together in a learned feature space while separating images from different classes. After this transformation, classification is performed using a linear decision boundary in that space. This ability to internally build such class-specific representations explains why classification accuracy remains higher for the visible neurons.
Overall, the high performance of the visible representations for both MNIST and CIFAR-10 demonstrates that the recalled representations remain highly class-discriminative and that the associative memory preserves the essential structure of the data, enabling compositional generalization to unseen examples.(for the CIFAR10 dataset, the classification accuracy can be increased by scaling up the associative memory, including $N_{h}$, the epoch size, the optimization step, and the number of training samples.)
@@
Table 1: Classification test accuracy of nonlinear classifiers trained and tested on recalled hidden and visible representations, as well as on the original images for comparison, for MNIST and CIFAR-10 datasets. High accuracy on visible representations for both datasets demonstrates that the recalled representations remain highly class-discriminative, while the lower accuracy on hidden representations for CIFAR-10 reflects the lack of strong class structure in raw pixel space.
4 Conclusion
This work introduces a novel Dense Associative Memory [krotov2021large] that achieves exponential storage capacity in the number of hidden neurons, overcoming the limitations of previous two-layer models. By using a threshold activation function, with a theoretically derived threshold, the network supports distributed hidden representations, allowing each hidden neuron to participate in multiple memories. This enables compositional storage of complex and correlated patterns, reducing redundancy while maintaining robust retrieval.
Specifically, the network achieves exponential capacity, $2^{N_{h}}$, using only $N_{h}N_{v}$ parameters. In contrast, previous two-layer implementations were limited to a maximum capacity of $N_{h}$ ([krotov2021large]). As a result, the number of memories per weight grows as $\frac{2^{N_{h}}}{N_{h}N_{v}}\approx 2^{N_{h}}$, while in previous implementations it is at best $\frac{1}{N_{v}}$. Even for complex datasets such as MNIST and CIFAR-10, networks with only 50 and 500 hidden units, respectively, were able to store tens of thousands of highly correlated memories and associate the vast majority of them with unique minima, whereas previous models could not store more memories than the number of hidden units (see AppendixC).
Beyond storing exponentially many memories, the network is also able to generalize to novel inputs. This behavior arises because the hidden layer encodes a set of basic memories that can be flexibly composed to represent previously unseen patterns, associating them with distinct minima rather than forcing retrieval toward the nearest stored pattern, while still producing meaningful, class-consistent representations. Our results are consistent with biological principles of feature learning, as embodied in hierarchical predictive coding models, which detect novelty and generalize by recombining features across successive levels of abstraction ([li2025predictive, salvatori2021associative]). This mechanism further illustrates that learning the underlying compositional structure of naturalistic data enables a biological associative memory to effectively support both memorization and generalization.
@@ Acknowledgments
We thank Brendan A. Bicknell and Reidar Riveland for helpful discussions. We acknowledge the following funding sources: Gatsby Charitable Foundation (GAT3850) to MSK and PEL.
@@ A.1 Distributional Properties of $\zeta_{\mu\nu}$
We define the matrix elements
$\zeta_{\mu\nu}=\frac{1}{\sqrt{N_{v}}}\sum_{i=1}^{N_{v}}{{\xi}}_{\mu i}{{\xi}}_{i\nu},\quad\mu\neq\nu,$ (18)
where ${{\xi}}_{\mu i}$s are randomly drawn from a standard normal distribution.
Each product ${{\xi}}_{\mu i}{{\xi}}_{i\nu}$ is a zero-mean random variable, since ${{\xi}}_{\mu i}$ and ${{\xi}}_{i\nu}$ are independent with zero mean.
By the central limit theorem, the sum of these $N_{v}$ independent terms converges in distribution to a Gaussian. Specifically,
$\sqrt{N_{v}}\left(\frac{1}{N_{v}}\sum_{i=1}^{N_{v}}{{\xi}}_{\mu i}{{\xi}}_{i\nu}-\mathbb{E}[{{\xi}}_{\mu i}{{\xi}}_{i\nu}]\right)\overset{d}{\longrightarrow}\mathcal{N}(0,1),$ (19)
given that $\mathbb{E}[{{\xi}}_{\mu i}{{\xi}}_{i\nu}]=0,$ and $\text{Var}[{{\xi}}_{\mu i}{{\xi}}_{i\nu}]=1$,
@@
Now consider the random variable $q_{\mu}$ defined as:
@@
$\text{Var}\left[\frac{1}{\sqrt{N_{v}}}\sum_{\nu=1}^{N_{h}}\zeta_{\mu\nu}s_{\nu}\right]=\frac{1}{N_{v}}\sum_{\nu=1}^{N_{h}}s_{\nu}^{2}\text{Var}[\zeta_{\mu\nu}]=\frac{1}{N_{v}}\sum_{\nu=1}^{N_{h}}s_{\nu}^{2}$ (22)
where we used the fact that the $\zeta_{\mu\nu}$ are independent random variables with mean 0 and variance 1.
@@ A.2 Stability of the fixed points
The stability of fixed points is determined by the Jacobian of the system. Grouping the variables into $(\mathbf{v},\mathbf{h})$, corresponding to the visible and hidden units respectively, the Jacobian has the block structure
$\mathbf{A}=\begin{bmatrix}\mathbf{A}_{vv}&\mathbf{A}_{vh}\\ \mathbf{A}_{hv}&\mathbf{A}_{hh}\end{bmatrix}$. (23)
@@ Appendix C Comparison with previous Dense Associative Memory networks
[krotov2021large] proposed a two-layer associative memory defined as
@@
where $w_{\mu i}=w_{i\mu}$ and each $\mathbf{w}_{\mu}\in\mathbb{R}^{N_{v}}$ represents a stored memory. Three choices for the nonlinearities $f$ and $g$ were introduced in [krotov2021large], summarized in Table 4.
We evaluate their recall performance (Models A, B, and C) together with our proposed model. Model A is tested using $f(h_{\mu})=h_{\mu}^{5}$ (A i) and $f(h_{\mu})=h_{\mu}^{10}$ (A ii).
@@
In contrast, our nonlinearity produces a fundamentally different recall regime. Multiple hidden neurons remain active simultaneously, enabling each hidden neuron to encode multiple stored patterns. This results in a dramatic increase in storage capacity: with only fifty hidden neurons, our model successfully stores all MNIST images (60,000) with high recall accuracy. By comparison, Models A, B, and C are limited to 50 memories and often fail to recall reliably (e.g., Model A i and Model C), as shown in Table 4.
@@ C $h_{\mu}^{5}$ $\dfrac{v_{i}}{\sqrt{\sum_{j}v_{j}^{2}}}$ 50 50 2%
Table 4: Nonlinearities used in the Dense Associative Memory models from [krotov2021large] and in our model, and a comparison of their recall performance. Recall performance is the percentage of recalled digits that are classified correctly.
Appendix D The ratio between the number of visible neurons and hidden neurons
From Eqs. (10) and (14), and setting $\theta=0.5$, we have
@@ Appendix D The ratio between the number of visible neurons and hidden neurons
where $q_{\mu,\text{target}}$ is a normally distributed random variable with variance $\frac{N_{a}}{N_{v}}$ (see 22), where $N_{a}$ is the number of active hidden neurons. The third term in the parentheses has variance $\sigma_{v}^{2}N_{h}/N_{v}$. Thus, in the limit $\tau_{v}\gg\tau_{h}\gg t$, we can approximate
@@
where
$z_{\mu,\text{target}}\sim\mathcal{N}\big(0,\sigma_{z}^{2}\big),\quad\sigma_{z}^{2}=\frac{N_{a}+\sigma_{v}^{2}N_{h}}{N_{v}}.$ (33)
Consequently, the probability of having no bit flip is
$P_{\text{no flip}}=\Phi(1/2\sigma_{z})^{N_{h}}$ (34)
where $\Phi$ is the cumulative normal function. This probability rapidly approaches 1 as the $\sigma_{z}$ becomes small. In this regime, both the stability of the fixed point and correct recall are ensured.
@@
And, to determine how many hidden neurons are required for real world memories with diverse statistics, note that memories from an $N_{v}$-dimensional space are recalled within an at most $N_{h}$-dimensional subspace spanned by the $N_{h}$ basic memories defined by the hidden to visible weights. If this subspace is not expressive enough, the reconstructed images will not be recognizable, particularly for complex datasets such as CIFAR-10.
In summary, the number of hidden neurons must be sufficiently smaller than the number of visible neurons to guarantee stable recall, but it must also be large enough to represent the statistical structure of the stored memories so that the reconstructions remain recognizable.
@@ Appendix F Sigmoid Activation Functions
We demonstrated this experimentally by showing that recall remains perfect even when a smooth sigmoid is used in place of the Heaviside function, as shown in Figure 10.
@@ Appendix G The Time Constant of Hidden and Visible Neurons
Figure 11 shows the impact of the time constant ratio between visible and hidden neurons on recall performance. As discussed in Section 3.2 on the basin of attraction, the visible neurons must be sufficiently slower than the hidden neurons because, during the cue, only the visible neurons receive input while the hidden neurons are initialized to zero. The visible neurons therefore need to evolve slowly enough to allow the hidden neurons to reach their correct steady state before the visible pattern changes significantly.

### v2

@@ 3.2 Experiments
Figure 3a shows the learned basic memories for the MNIST dataset, which correspond to the columns of $\xi$}. As shown in Figure 3b, these basic memories are nearly orthogonal, consistent with Eq. (7a).
Figure 3: a) 25 (out of 50) columns of the learned weight matrix for MNIST images, which serve as basic memories, are shown as two-dimensional images. b) Correlation matrix of the basic patterns, which correspond to the hidden units. c) The network generalizes compositionally, associating unseen cues with interpretable fixed points.
@@
The network learns a global threshold of $\theta=0.21$ for MNIST and $\theta=0.43$ for CIFAR-10. Note that the statistics of the learned basic memories in MNIST and CIFAR-10 differ from one another and from the normal distribution assumed in the theory, which explains the difference between the learned thresholds.
@@
To quantitatively evaluate this behavior, we trained nonlinear classifiers on the recalled hidden and visible representations of the stored images, and tested them on the recalled representations of unseen images (Convolutional neural network (CNN) for visible representations and multilayer perceptron (MLP) for hidden representations). For comparison, we trained CNNs directly on the original stored images and tested them on the original unseen images as well. This allows us to assess the classifiability of the recalled visible representations by the associative memory network relative to the original memories. (see Appendix B for details of the classifiers)
For MNIST, classification accuracy is high for both the hidden and visible representations. This is a desirable property, as it indicates that the lower-dimensional hidden representations still preserve strong class discriminability. In MNIST, the raw pixel space itself carries strong class structure: two images of the same digit are highly correlated and closer to each other than images of different digits. Consequently, the hidden neurons retain this information almost perfectly, having structured and meaningful encodings in which correlated memories are represented close together and remain classifiable.
For CIFAR-10, classification accuracy is higher for the visible representations than for the hidden ones. This difference arises because, in CIFAR-10, two images of the same class (for example, dogs) are not necessarily correlated in raw pixel space, so the hidden layer, which is a nonlinear transformation of those pixels, does not exhibit a clear class structure. The classifier used for the visible representations in this analysis is a CNN, which, when trained on raw pixel data, first learns a nonlinear transformation that maps images of the same class close together in a learned feature space while separating images from different classes. After this transformation, classification is performed using a linear decision boundary in that space. This ability to internally build such class-specific representations explains why classification accuracy remains higher for the visible neurons.
Overall, the high performance on the visible representations for MNIST, and the reasonably high performance for CIFAR-10, demonstrates that the recalled representations remain class-discriminative, and that the associative memory preserves the essential structure of the data, enabling compositional generalization to unseen examples (for the CIFAR10 dataset, the classification accuracy can be increased by scaling up the associative memory, including $N_{h}$, the epoch size, the optimization step, and the number of training samples).
@@
Table 1: Classification test accuracy of nonlinear classifiers trained and tested on recalled hidden and visible representations, as well as on the original images for comparison, for MNIST and CIFAR-10. Reasonably high accuracy on visible representations for both datasets demonstrates that the recalled representations remain highly class-discriminative, while the lower accuracy on hidden representations for CIFAR-10 reflects the lack of strong class structure in raw pixel space.
4 Conclusions
This work introduces a novel Dense Associative Memory Krotov and Hopfield (2021) that achieves exponential storage capacity in the number of hidden neurons, overcoming the limitations of previous two-layer models. By using a threshold activation function, with a theoretically derived threshold, the network supports distributed hidden representations, allowing each hidden neuron to participate in multiple memories. This enables compositional storage of complex and correlated patterns, reducing redundancy while maintaining robust retrieval.
Notably, the network achieves exponential capacity, $2^{N_{h}}$, using only $N_{h}N_{v}$ parameters. In contrast, previous two-layer implementations were limited to a maximum capacity of $N_{h}$ (Krotov and Hopfield (2021)). As a result, in the proposed network, the number of memories per weight grows as $\frac{2^{N_{h}}}{N_{h}N_{v}}\approx 2^{N_{h}}$, while in previous implementations it is at best $\frac{1}{N_{v}}$. Even for complex datasets such as MNIST and CIFAR-10, networks with only 50 and 500 hidden units, respectively, were able to store tens of thousands of highly correlated memories and associate the vast majority of them with unique minima, whereas previous models could not store more memories than the number of hidden units (see AppendixC).
Beyond storing exponentially many memories, the network is also able to generalize to novel inputs. This behavior arises because the hidden layer encodes a set of basic memories that can be flexibly composed to represent previously unseen patterns, associating them with distinct minima rather than forcing retrieval toward the nearest stored pattern, while still producing meaningful, class-consistent representations. Our results are consistent with biological principles of feature learning, as embodied in hierarchical predictive coding models, which detect novelty and generalize by recombining features across successive levels of abstraction (Li et al. (2025); Salvatori et al. (2021)). This mechanism further illustrates that learning the underlying compositional structure of naturalistic data enables a biological associative memory to effectively support both memorization and generalization.
@@ Acknowledgments
We thank Brendan A. Bicknell and Reidar Riveland for helpful discussions. We acknowledge the following funding sources: Gatsby Charitable Foundation (GAT3850) to MSK and PEL. The results presented here were obtained while Dmitry Krotov was employed by IBM Research. At the time of the camera-ready submission Dmitry Krotov is no longer employed by IBM Research.
@@ A.1 Distributional Properties of $\zeta_{\mu\nu}$
Combining Eqs. (6) and (7a), we see that $\zeta_{\mu\nu}$ is given by
$\zeta_{\mu\nu}=\frac{1}{\sqrt{N_{v}}}\sum_{i=1}^{N_{v}}{{\xi}}_{\mu i}{{\xi}}_{i\nu}-\sqrt{N_{v}}\delta_{\mu\nu}\,.$ (18)
where, recall, the ${{\xi}}_{\mu i}$ are independent, zero mean unit variance Gaussian random variables (see Eq. (3)). Thus, for $\mu\neq\nu$, each product ${{\xi}}_{\mu i}{{\xi}}_{i\nu}$ is a zero-mean, unit variance random variable. By the central limit theorem, the sum of these $N_{v}$ independent terms converges in distribution to a Gaussian, so for $\mu\neq\nu$ we have
@@
For $\mu=\nu$, we write
$\zeta_{\mu\mu}=\frac{1}{\sqrt{N_{v}}}\sum_{i=1}^{N_{v}}({{\xi}}_{\mu i}^{2}-1)\,.$ (20)
By the central limit theorem, this too is Gaussian, but with a slightly larger variance,
$\zeta_{\mu\mu}\overset{d}{\longrightarrow}\mathcal{N}(0,\textrm{Var}[{{\xi}}^{2}])=\mathcal{N}(0,2)\,.$ (21)
Combining Eqs. (19) and (21), we arrive at
$\zeta_{\mu\nu}\overset{d}{\longrightarrow}\mathcal{N}(0,1+\delta_{\mu\nu})\,.$ (22)
Now consider the random variable $q_{\mu}$ defined as
@@
$\text{Var}\left[\frac{1}{\sqrt{N_{v}}}\sum_{\nu=1}^{N_{h}}\zeta_{\mu\nu}s_{\nu}\right]=\frac{1}{N_{v}}\sum_{\nu=1}^{N_{h}}s_{\nu}^{2}\text{Var}[\zeta_{\mu\nu}]=\frac{1}{N_{v}}\sum_{\nu=1}^{N_{h}}s_{\nu}^{2}(1+\delta_{\mu\nu})$ (24)
where we used the fact that the $\zeta_{\mu\nu}$ are independent random variables with mean 0 and variance given in Eq. (22).
@@ A.2 Stability of the fixed points
The stability of fixed points is determined by the Jacobian of the system. Grouping the variables into $(\mathbf{v},\mathbf{h})$, corresponding to the visible and hidden units respectively, the Jacobian, denoted $\mathbf{A}$, has the block structure
$\mathbf{A}=\begin{bmatrix}\mathbf{A}_{vv}&\mathbf{A}_{vh}\\ \mathbf{A}_{hv}&\mathbf{A}_{hh}\,.\end{bmatrix}$ (25)
@@ Appendix C Comparison with previous Dense Associative Memory networks
Krotov and Hopfield [2021] proposed a two-layer associative memory defined as
@@
where $w_{\mu i}=w_{i\mu}$ and each $\mathbf{w}_{\mu}\in\mathbb{R}^{N_{v}}$ represents a stored memory. Three choices for the nonlinearities $f$ and $g$ were introduced in Krotov and Hopfield [2021], summarized in Table 4.
We evaluate their recall performance (Models A, B, and C) together with our proposed model. Model A is tested using $f(h_{\mu})=h_{\mu}^{5}$ (A (i)) and $f(h_{\mu})=h_{\mu}^{10}$ (A (ii)).
@@
In contrast, our nonlinearity produces a fundamentally different recall regime. Multiple hidden neurons remain active simultaneously, enabling each hidden neuron to encode multiple stored patterns. This results in a dramatic increase in storage capacity: with only fifty hidden neurons, our model successfully stores all MNIST images (60,000) with high recall accuracy. By comparison, Models A, B, and C are limited to 50 memories and often fail to recall reliably (e.g., Model A (i) and Model C), as shown in Table 4.
@@ C $h_{\mu}^{5}$ $\dfrac{v_{i}}{\sqrt{\sum_{j}v_{j}^{2}}}$ 50 50 2%
Table 4: Nonlinearities used in the Dense Associative Memory models from Krotov and Hopfield [2021] and in our model, and a comparison of their recall performance. Recall performance is the percentage of recalled digits that are classified correctly.
Appendix D Noise-induced bit flips
As pointed out in Secs. 2.1 and 2.2, if the noise – associated either with finite $N_{v}$ corrections to $J_{\mu\nu}$, with perturbations to the initial value of the visible units, or both – is too large, the target solution can be destabilized via “bit flips” in the hidden units. Here we quantify this, and in particular show that the probability of a bit flip falls off rapidly as the number of visible units, $N_{v}$, increases.
From Eqs. (9) and (14), with $\theta$ set to 0.5, we have
@@ Appendix D Noise-induced bit flips
where $q_{\mu,\text{target}}$ is a normally distributed random variable with variance given in Eq. (24), but with $s_{\nu}$ replaced by $s_{\nu,\text{target}}$. As discussed in the paragraph following Eq. (14), the third term in parentheses has variance $\sigma_{v}^{2}N_{h}/N_{v}$. Thus, in the limit $\tau_{v}\gg\tau_{h}\gg t$, we can approximate
@@
where $z_{\mu,\text{target}}$ is normally distributed with variance given by, at most,
$\text{Var}[z_{\mu,\text{target}}]=\frac{N_{a,\text{target}}+1+\sigma_{v}^{2}N_{h}}{N_{v}}\,.$ (35)
Here $N_{a,\text{target}}$ is the number of active target hidden neurons, and the “at most” qualifier is because we should include the “+1” term only if $s_{\mu,\text{target}}=1$ (see Eq. (24)).
The probability of a single bit flip depends in a relatively simple way on the variance of $z_{\mu,\text{target}}$. However, the dependence on the number of active target neurons and the “+1” makes it nontrivial to determine the probability that there are no bit flips at all. We therefore bound that probability by focusing on the upper bound on the variance of $z_{\mu,\text{target}}$, which is given by
$\sigma_{z}^{2}\equiv\frac{N_{h}+1+\sigma_{v}^{2}N_{h}}{N_{v}}\,.$ (36)
With this definition, the probably of a single bit flip , which is the probability that $z_{\mu,\text{target}}$ is either greater than 1/2 or less than -1/2, depending on whether $s_{\mu,\text{target}}$ is 0 or 1, is bounded by
$P_{\text{1 bit flip}}\leq 1-\Phi(1/2\sigma_{z})$ (37)
where $\Phi$ is the cumulative normal function. The probability of no bit flips is, then, bounded by
$P_{\text{no bit flips}}=(1-P_{\text{1 bit flip}})^{N_{h}}\geq\Phi(1/2\sigma_{z})^{N_{h}}\,.$ (38)
To write this in a more interpretable form, note that
$1-\Phi(a)=\int_{a}^{\infty}dx\,\frac{e^{-\frac{x^{2}}{2}}}{\sqrt{2\pi}}=\int_{0}^{\infty}dy\,\frac{e^{-\frac{a^{2}}{2}-ay-\frac{y^{2}}{2}}}{\sqrt{2\pi}}\leq\int_{0}^{\infty}dy\,\frac{e^{-\frac{a^{2}}{2}-ay}}{\sqrt{2\pi}}=\frac{e^{-\frac{a^{2}}{2}}}{a\sqrt{2\pi}}\,.$ (39)
Inserting this into Eq. (40) then gives us the bound
$P_{\text{no bit flips}}\geq\left(1-2\sigma_{z}\,\frac{e^{-1/8\sigma_{z}^{2}}}{\sqrt{2\pi}}\right)^{N_{h}}\geq 1-N_{h}\sigma_{z}\,\frac{e^{-1/8\sigma_{z}^{2}}}{\sqrt{\pi/2}}$ (40)
where the second inequality follows from Jensen’s inequality.
This probability rapidly approaches 1 as $\sigma_{z}$ becomes small. In this regime, both the stability of the fixed point and correct recall are ensured.
@@
To determine how many hidden neurons are required for real world memories with diverse statistics, note that memories from an $N_{v}$-dimensional space are recalled within an at most $N_{h}$-dimensional subspace spanned by the $N_{h}$ basic memories defined by the hidden to visible weights. If this subspace is not expressive enough, the reconstructed images will not be recognizable, particularly for complex datasets such as CIFAR-10. Consequently, the number of hidden neurons must be sufficiently smaller than the number of visible neurons to guarantee stable recall, but it must also be large enough to represent the statistical structure of the stored memories so that the reconstructions remain recognizable.
@@ Appendix F Sigmoid Activation Functions
We demonstrated this experimentally by showing that recall remains perfect even when a smooth sigmoid is used in place of the Heaviside function, as shown in Figure 10b.
@@ Appendix G The Time Constant of Hidden and Visible Neurons
Figure 11 shows the impact of $\tau_{v}/\tau_{h}$, the ratio of the time constant between visible and hidden neurons on recall performance. As discussed in Section 3.2 on the basin of attraction, the visible neurons must be sufficiently slower than the hidden neurons because, during the cue, only the visible neurons receive input while the hidden neurons are initialized to zero. The visible neurons therefore need to evolve slowly enough to allow the hidden neurons to reach their correct steady state before the visible pattern changes significantly. Somewhat surprisingly, a factor of only 4 is enough to ensure perfect recall.